OpenStudy (anonymous):

verify 1+sinΘ/cosΘ + cosΘ sinΘ = secΘ

OpenStudy (abb0t):

combine like terms.

OpenStudy (anonymous):

ok

OpenStudy (abb0t):

\[\frac{1+\sin(\theta) + \cos(\theta)\cos(\theta)\sin(\theta) }{ \cos(\theta) }=...\]

OpenStudy (anonymous):